# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v3

@@ Appendix C More Results: Evaluation with Proprietary LLM
Given the resource-intensive nature of LLM fine-tuning, we conducted further experiments in coherence evaluation using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o. These advanced and widely used models offer strong performance in various tasks but also come with higher operational costs due to token consumption. Using the same evaluation set of 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively. The results highlight that a smaller, fine-tuned LLM, such as Qwen-2.5-7b (with an accuracy of 96.25%), can outperform these advanced models while significantly reducing costs associated with API usage [^4] The fine-tuned Chinese LLMs for coherence evaluation and cluster naming will be released upon acceptance.. For comparison, a vanilla Qwen-2.5-7b model (without fine-tuning) was also evaluated, yielding a substantially lower accuracy of 75.63%, emphasizing the essence of fine-tuning.
{CJK}
UTF8gkai
Cluster Coherence Original Sentences English Translation
Good "给企业固定资产买保险,大约能投大约得投保多少呢", "就是假如我有一百万的企业固定资金买保险大概投保多少钱啊", "请问一下我想咨>询一下企业财产保险", "您好我想了解一下这个企业财产保险", "就了解一下这个企业财产保险","你们是财险", "是哦这个属于财险了对吧", "企业财产保险的", "这些都属于财产险对吗", "财险人工那你这是", 就是如果我要为我这个私营企业买这个保险需要什么手续", "财产险", "呃企业财产保险是以什么为保" "If I buy insurance for a company’s fixed assets, how much insurance will it cost?", "That is, if I have a million corporate fixed assets, how much will it cost to buy insurance?", "Excuse me, I would like to consult> Ask about corporate property insurance", "Hello, I want to know about this corporate property insurance", "Just want to know about this corporate property insurance", "You are a property insurance company", "Yes, this belongs to property insurance "Right?", "Enterprise property insurance", "These all belong to property insurance, right?", "What about property insurance workers?", What do I need if I want to buy this insurance for my private enterprise? Procedure","Property Insurance", "Well, what does corporate property insurance cover?"
Bad "就是连连续,就是一直一直保", "但是它连不上是怎么回事啊?", "哦就是主要是直接给公司转账对吧?", "接吗", "接也是吗?", "我直>接去", "你是直接直接用那个" , "就是从哪接过来再接回去","还是需要从哪儿连线这个宽带", "你直接给我说这些啊", "直接把", "哦直接", "嗯那个礼品是直接就发放了", "是你直接给我回复对吗", "直接到那里去", "嗯,最好直飞。", "请帮我连接+", "把钱打过去的话,我是直接打到那个证券公司", "直接就是" "It’s continuous, it’s always guaranteed", "But what’s wrong with it not being able to connect?", "Oh, it’s mainly to transfer money directly to the company, right?", "Yes", "Yes too?" ?", "I’ll go directly", "Are you using that directly" , "Just connect it from where you are and then connect it back", "Or do you need to connect to the broadband from somewhere", "You Just tell me this directly", "Just give it directly", "Oh directly", "Well, the gift was given out directly", "You replied to me directly, right?","Go there directly" , "Well, it’s best to fly directly.", "Please help me connect +", "If I call the money, I will call the securities company directly", "Directly"
@@
{CJK}
UTF8gkai
Cluster Name Original Sentences English Translation
询问-优惠 (Inquire-Promotion) ’那有什么优惠券什么之类的吗?’, ’是是怎么形式是优惠吗?’, ’还是不是优惠活动,是那个直接给我打我卡里吗 ,还是是什么优惠券儿啊?’, ’就比如新用户他有什么优惠券儿之类的吧!’, ’嗯你你那还有什么优惠的活动吗? 就是比较合适就是合适的动。’ ’Are there any coupons or anything like that? ’, ’What is the form of the discount? ’, ’Still, it’s not a promotion. Is it the one that directly charges my card? , or is it some kind of coupon? ’, ’For example, what coupons does a new user have? ’, ’Well, do you have any other discounts? It is more appropriate and appropriate to move. ’
解答-金额 (Answer-Amount) ’一共是三十一块二毛’, ’就每个月一百三十八’, ’对,一个月也就是四百百四五百块钱嘛, 给您自己做个积累。’, ’十二月份的话是用了三十三块九毛二’, ’对一个月一百三十八’ ’The total is thirty-one and twenty cents’, ’That’s one hundred and thirty-eight cents per month’, ’Yes, that’s four hundred, four hundred and five hundred yuan per month. Make an accumulation for yourself. ’, ’In December, it cost thirty-three dollars and ninety-two cents’, ’That’s one hundred and thirty-eight dollars a month’
@@
Table 10: Example of cluster naming with "Action-Objective" style.
Dataset Naming Convention Example (sentence) Example (label)
NLU scenario-intent Send an email to Alex and write thank you. email sendemail
NLU++ list of keywords How long does it usually take to get a new pin? ["how_long","pin", "arrival","new"]
OOS objective Please tell me why my card was declined yesterday. card_declined
ours action-objective Well, do you have any other discounts? inquire-promotion
@@
Table 11: Comparison of cluster naming convention in intent clustering datasets.
{CJK}
UTF8gkai
Cluster Name Clusters with Similar Intention English Translation
询问-意外事故 (Inquire-Accident) "假如被车碰了或者是被楼上的砖砸了一下", "给别人儿撞的意外", "啊撞到别人然后就是", "就平时有时候开车嘛可能会遇到这个", "嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后", "就是把其他的东西撞到了呀什么的", "那人生意,不是就是,如果是不小心在马上被车撞了的话", "撞到人了吧", "然后不小在行驶当中被别人损害就是说拿石头砸的呀然后", "什么被车撞了之类的,是吗", "被被被撞了,还是被什么一些什么意外事故了", "我把别人的车撞了是吧" "If I were hit by a car or hit by a brick from upstairs," "Accidentally hit by someone else," "Ah, hit someone else and then," "Just sometimes when driving, you might encounter this," "Well, if I’m not careful and hit some place myself," "Just hit something else or something," "That’s a human affair, not just, if it’s accidentally hit by a car on the road," "Hit someone, right?" "Then not small in the process of driving, being damaged by someone else, say, hit with a stone, and then," "What, hit by a car or something, right?" "I hit someone else’s car, right?"
询问-意外死亡 (Inquire-Accident Death) "哦猝死,那猝死算意外吗", "啊,那我想知道那个猝死的话,算是意外死亡吗", "算意外死亡吗", "那如果猝死算是意外死亡吗?", "猝死也算意外事吧", "那那猝死是意外死亡吗如果是猝死的话是", "猝死算是意外死亡吗", "猝死算意外死亡吗", "嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗", "那猝死的话,算意外死亡吗?", "那那那那个就是那个猝死算是意外死亡吗" "Oh sudden death, is sudden death considered accidental?" "Ah, I want to know if sudden death is considered accidental death?" "Is it considered accidental death?" "If sudden death is considered accidental death?" "Sudden death is also an accident, right?" "Is sudden death considered accidental death if it is sudden death?" "Is sudden death considered accidental death?" "Is sudden death considered accidental death?" "Well, that sudden death, does sudden death fall under accidental insurance, accidental death?" "Is sudden death considered accidental death?" "Is that, that, that sudden death considered accidental death?"
@@
Table 13: Example of log record for iterative intent clustering.

### v4

@@ Appendix B Prompt Template
Model Coherence Evaluation Cluster Naming
LLaMA-2-7B (LoRA) 94.0% $\pm$ 1.2% 89.0% $\pm$ 1.5%
LLaMA-2-13B (LoRA) 95.0% $\pm$ 1.1% 93.5% $\pm$ 1.4%
Mistral-7B (LoRA) 91.5% $\pm$ 2.3% 82.0% $\pm$ 1.9%
GPT-3.5 (API) 91.0% $\pm$ 1.5% 86.0% $\pm$ 1.8%
GPT-4 (API) 94.5% $\pm$ 1.2% 91.0% $\pm$ 1.3%
Table 10: Performance of different LLMs on coherence evaluation and cluster naming for the English dataset.
@@ Appendix C More Results: Evaluation with Proprietary LLM
Given the resource-intensive nature of LLM fine-tuning, we conducted additional experiments using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o. These widely adopted models deliver strong performance across diverse tasks and do not require task-specific fine-tuning, thereby alleviating the data scarcity issue, but they also incur higher operational costs due to token consumption. On the same Chinese coherence evaluation dataset with 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively. Notably, the smaller fine-tuned Qwen-2.5-7B reached 96.25% accuracy, surpassing these advanced proprietary models while significantly reducing API-related costs. For comparison, the vanilla Qwen-2.5-7B model (without fine-tuning) obtained a much lower accuracy of 75.63%, further underscoring the importance of fine-tuning.
On the English dataset, we report results for GPT-3.5 and GPT-4, as well as two additional fine-tuned models: LLaMA-2-13B (LoRA) and Mistral-7B (LoRA). The LLaMA and Mistral models were fine-tuned on the same 800 intent clusters as LLaMA-2-7B in Section 5.3 and evaluated on 100 English intent clusters annotated by human experts with Good/Bad and Action-Objective intent labels. Each model was run five times with different random seeds to ensure robust performance metrics. Results in Table 10 show that scaling up model size (e.g., 7B vs. 13B) improves performance, while fine-tuned smaller LLMs often outperform large proprietary LLMs, consistent with the findings on the Chinese dataset. These results reinforce the value of fine-tuning and suggest that smaller, cost-efficient models can play a critical role in data mining within the LLM-in-the-loop framework.
Cluster Coherence
Original Sentences
English Translation
Good
”给企业固定资产买保险,大约能投大约得投保多少呢”,
”就是假如我有一百万的企业固定资金买保险大概投保多少钱啊”,
”请问一下我想咨¿询一下企业财产保险”,
”您好我想了解一下这个企业财产保险”,
”就了解一下这个企业财产保险”,”你们是财险”,
”是哦这个属于财险了对吧”, ”企业财产保险的”,
”这些都属于财产险对吗”, ”财险人工那你这是”,
就是如果我要为我这个私营企业买这个保险需要什么手续”,
”财产险”, ”呃企业财产保险是以什么为保”
”If I buy insurance for a company’s fixed assets,
how much insurance will it cost?”,
”That is, if I have a million corporate fixed assets,
how much will it cost to buy insurance?”,
”Excuse me, I would like to consult¿ Ask about corporate property insurance”,
”Hello, I want to know about this corporate property insurance”,
”Just want to know about this corporate property insurance”,
”You are a property insurance company”,
”Yes, this belongs to property insurance ”Right?”, ”Enterprise property insurance”,
”These all belong to property insurance, right?”,
”What about property insurance workers?”,
What do I need if I want to buy this
insurance for my private enterprise? Procedure”,”Property Insurance”,
”Well, what does corporate property insurance cover?”
Bad
”就是连连续,就是一直一直保”, ”但是它连不上是怎么回事啊?”,
”哦就是主要是直接给公司转账对吧?”, ”接吗”, ”接也是吗?”,
”我直¿接去”, ”你是直接直接用那个”
, ”就是从哪接过来再接回去”,”还是需要从哪儿连线这个宽带”,
”你直接给我说这些啊”,
”直接把”, ”哦直接”, ”嗯那个礼品是直接就发放了”,
”是你直接给我回复对吗”, ”直接到那里去”,
”嗯,最好直飞。”, ”请帮我连接+”,
”把钱打过去的话,我是直接打到那个证券公司”, ”直接就是”
”It’s continuous, it’s always guaranteed”,
”But what’s wrong with it not being able to connect?”,
”Oh, it’s mainly to transfer money directly to the company, right?”,
”Yes”, ”Yes too?” ?”,
”I’$l$ go directly”, ”Are you using that directly”
, ”Just connect it from where you are and then connect it back”,
”Or do you need to connect to the broadband from somewhere”,
”You Just tell me this directly”, ”Just give it directly”, ”Oh directly”,
”Well, the gift was given out directly”,
”You replied to me directly, right?”,”Go there directly” ,
”Well, it’s best to fly directly.”, ”Please help me connect +”,
”If I call the money, I will call the securities company directly”, ”Directly”
@@
Cluster Name
Original Sentences
English Translation
询问-优惠
(Inquire-Promotion)
’那有什么优惠券什么之类的吗?’,
’是是怎么形式是优惠吗?’,
’还是不是优惠活动,是那个直接给我打我卡里吗
,还是是什么优惠券儿啊?’,
’就比如新用户他有什么优惠券儿之类的吧!’,
’嗯你你那还有什么优惠的活动吗?
就是比较合适就是合适的动。’
’Are there any coupons or anything like that? ’,
’What is the form of the discount? ’,
’Still, it’s not a promotion. Is it the one that directly charges my card?
, or is it some kind of coupon? ’,
’For example, what coupons does a new user have? ’,
’Well, do you have any other discounts?
It is more appropriate and appropriate to move. ’
解答-金额
(Answer-Amount)
’一共是三十一块二毛’,
’就每个月一百三十八’,
’对,一个月也就是四百百四五百块钱嘛,
给您自己做个积累。’,
’十二月份的话是用了三十三块九毛二’,
’对一个月一百三十八’
’The total is thirty-one and twenty cents’,
’That’s one hundred and thirty-eight cents per month’,
’Yes, that’s four hundred,
four hundred and five hundred yuan per month. Make an accumulation for yourself. ’,
’In December, it cost thirty-three dollars and ninety-two cents’,
’That’s one hundred and thirty-eight dollars a month’
@@
Table 12: Example of cluster naming with ”Action-Objective” convention.
Dataset
Naming Convention
Example (sentence)
Example (label)
NLU
scenario-intent
Send an email to Alex and write thank you.
email sendemail
NLU++
list of keywords
How long does it usually take to get a new pin?
[”how_long”,”pin”,
”arrival”,”new”]
OOS
objective
Please tell me why my card was declined yesterday.
card_declined
ours
action-objective
Well, do you have any other discounts?
inquire-promotion
@@
Table 13: Comparison of cluster naming conventions in existing and proposed intent clustering datasets.
Cluster Name
Clusters with Similar Intention
English Translation
询问-意外事故
(Inquire-Accident)
”假如被车碰了或者是被楼上的砖砸了一下”,
”给别人儿撞的意外”,
”啊撞到别人然后就是”,
”就平时有时候开车嘛可能会遇到这个”,
”嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后”,
”就是把其他的东西撞到了呀什么的”,
”那人生意,不是就是,如果是不小心在马上被车撞了的话”,
”撞到人了吧”,
”然后不小在行驶当中被别人损害就是说拿石头砸的呀然后”,
”什么被车撞了之类的,是吗”,
”被被被撞了,还是被什么一些什么意外事故了”,
”我把别人的车撞了是吧”
”If I were hit by a car or hit by a brick from upstairs,”
”Accidentally hit by someone else,”
”Ah, hit someone else and then,”
”Just sometimes when driving, you might encounter this,”
”Well, if I’m not careful and hit some place myself,”
”Just hit something else or something,”
”That’s a human affair, not just, if it’s accidentally
hit by a car on the road,”
”Hit someone, right?”
”Then not small in the process of driving, being
damaged by someone else, say, hit with a stone, and then,”
”What, hit by a car or something, right?”
”I hit someone else’s car, right?”
询问-意外死亡
(Inquire-Accident Death)
”哦猝死,那猝死算意外吗”,
”啊,那我想知道那个猝死的话,算是意外死亡吗”,
”算意外死亡吗”,
”那如果猝死算是意外死亡吗?”,
”猝死也算意外事吧”,
”那那猝死是意外死亡吗如果是猝死的话是”,
”猝死算是意外死亡吗”,
”猝死算意外死亡吗”,
”嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗”,
”那猝死的话,算意外死亡吗?”,
”那那那那个就是那个猝死算是意外死亡吗”
”Oh sudden death, is sudden death considered accidental?”
”Ah, I want to know if sudden death is considered accidental death?”
”Is it considered accidental death?”
”If sudden death is considered accidental death?”
”Sudden death is also an accident, right?”
”Is sudden death considered accidental death if it is sudden death?”
”Is sudden death considered accidental death?”
”Is sudden death considered accidental death?”
”Well, that sudden death, does sudden death fall under
accidental insurance, accidental death?”
”Is sudden death considered accidental death?”
”Is that, that, that sudden death considered accidental death?”
@@
Model Sampling Method 1 2 3 4 5
qwen14b convex - 96.25% 95.42% 95.42% 95.21%
random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%
Table 15: Comparison of sampling methods and hyperparameters for LLM coherence evaluation.
@@
Table 16: Example log records from iterative intent clustering across epochs.